OpenStudy (anonymous):

Simplify. n ^6 · n^ 5 ÷ n ^4 · n ^3 ÷ n ^2 · n

OpenStudy (anonymous):

\[(n^6n^5)/(n^4n^3)/(n^2n)=n^{11}/n^7/n^3=n^{11}(n^3/n^7)=n^{14}/n^7\] =\[n^{14-7}= n^7\]

OpenStudy (anonymous):

remember\[a^ma^n=a^{m+n}, a^m/a^n=a^{m-n}\] basic exponent properties